Translate the statement into an equation using “K” as the constant of proportionality?
I is directly proportional to the cube root of y if I= 5 when y=64 find I when y=8

Answer by Boreal(15235)

I=ky^(1/3)
5=k*64^(1/3)=4k
k=5/4
use that in the second part
I=(5/4)*(8)^(1/3)
=5/2